OpenStudy (devante):

how do you simplify this expression 3x4y2/xy2?

OpenStudy (anonymous):

cancel terms that r common in num nd denom :P

OpenStudy (anonymous):

that how u simplify 3x3y2

OpenStudy (anonymous):

no y2 sry

OpenStudy (apoorvk):

Use the funda: \[1. ~a^m \times a^n = a^{m+n}\]and \[2. \frac{a^m}{a^n} = a^{m-n}\]